In reaching her destination, a backpacker walks with an average velocity of 1.17 m/s, due west. This average velocity results, because she hikes for 5.57 km with an average velocity of 3.24 m/s due west, turns around, and hikes with an average velocity of 0.486 m/s due east. How far east did she walk (in kilometers)?
Answer #1

Consider west as positive direction of motion and east as negative direction of motion

Vavg = average velocity = 1.17 m/s

Dw = distance traveled in west direction = 5.57 km = 5570 m

vw = average velocity to travel distance in west direction = 3.24 m/s

tw = time of travel in west direction = Dw /vw = 5570/3.24 = 1719.14 sec

DE = distance traveled in east direction = ?

vE = average velocity to travel distance in east direction = - 0.486 m/s

tE = time of travel in east direction = DE /vE

Average velocity is given as

Vavg = Total displacement/total time

Vavg = (Dw - DE ) /(tw + tE )

1.17 = (5570 - DE ) /(1719.14 + DE /0.486)

DE = 1044.4 m

DE = 1.044 km